### v1

# LSAQ: Layer-Specific Adaptive Quantization for Large Language Model Deployment

Abstract
As large language models (LLMs) demonstrate exceptional performance across various domains, the deployment of these models on edge devices has emerged as a new trend. Quantization techniques, which reduce the size and memory footprint of LLMs, are effective for enabling deployment on resource-constrained edge devices. However, existing one-size-fits-all quantization methods often fail to dynamically adjust the memory consumption of LLMs based on specific hardware characteristics and usage scenarios. To address this limitation, we propose LSAQ (Layer-Specific Adaptive Quantization), a system for adaptive quantization and dynamic deployment of LLMs based on layer importance. LSAQ evaluates layer importance by constructing top-$k$ token sets from the inputs and outputs of each layer and calculating their Jaccard coefficient. Using this evaluation, the system adaptively adjusts quantization strategies in real time according to the resource availability of edge devices, assigning different precision levels to layers of varying importance. This approach significantly reduces the storage requirements of LLMs while maintaining model performance, enabling efficient deployment across diverse hardware platforms and usage scenarios.
Index Terms:
large language models, quantization, layer importance, jaccard coefficient, edge devices
I Introduction
With the rapid advancement of large language models (LLMs), they have demonstrated exceptional capabilities across a variety of tasks. Consequently, the use of LLMs has proliferated across multiple domains, including natural language processing [1], code generation [2, 3], and even in finance [4] and education [5] sectors. To harness LLMs for more intelligent and personalized services [6] while ensuring data security and privacy, the deployment of LLMs on edge devices has emerged as a new trend. However, the massive parameter size of these models poses significant challenges for efficient deployment on edge devices. For example, a model with tens of billions of parameters, such as Llama-2-13B [7], requires approximately 25 GB of memory just to load its weights. In contrast, a high-performance consumer-grade GPU such as the NVIDIA GeForce RTX 4090 has only 24 GB of memory, rendering it incapable of deploying Llama-2-13B directly. To effectively address this challenge, researchers are exploring model compression techniques such as quantization [8, 9], pruning [10, 11, 12], distillation [13, 14], and low-rank factorization [15] to optimize the storage and computational requirements of LLMs.
In the field of model compression, model quantization plays a pivotal role. Its core mechanism involves converting high-precision floating-point weights or activation values in a model into lower-precision numerical representations. This transformation significantly reduces the size of the model storage and minimizes memory consumption during inference. Post-training quantization (PTQ), a widely adopted quantization technique, is particularly valuable for deploying models on resource-constrained devices, such as mobile devices and embedded systems. Quantization methods such as LLM.int8() [9], GPTQ [8], and Omniquant [16] fall under the PTQ category.
However, many prevalent PTQ methods uniformly quantize across all layers of LLMs, neglecting the varying significance of these layers. This practice disregards evidence from existing studies [21] that indicate substantial redundancy in numerous layers within LLMs. This one-size-fits-all quantization approach also limits the ability to dynamically adjust the memory consumption of LLMs based on specific hardware characteristics and task requirements. The concept of Layer-Wise Quantization [17] offers a potential solution to this problem by quantifying each layer based on the cosine similarity between the input and output of the layer, or the distribution of weight outliers, as indicators of layer importance. Nevertheless, these methods have limitations: neither directly reflects the degree of semantic change, and the latter struggles to effectively distinguish layers with relatively uniform weight distributions. Therefore, it is crucial to explore more refined quantization strategies to enable differentiated treatment of LLM layers. Additionally, techniques capable of dynamically adjusting the quantization precision based on practical application scenarios should be developed to further optimize the storage and inference efficiency while maintaining model performance.
Figure 1: Framework of LSAQ system
To that end, we propose LSAQ, a novel Layer-Specific Adaptive Quantization framework tailored for the deployment of LLMs, as illustrated in Figure 1. LSAQ introduces a fine-grained layer importance evaluation mechanism by constructing top-$k$ token sets corresponding to the inputs and outputs of each layer in LLMs. It calculates the Jaccard coefficient between these two sets as an indicator of layer importance. A higher Jaccard coefficient implies greater similarity between the input and output token sets, suggesting that the layer has not achieved significant semantic transformation when processing the input. Consequently, such layers are deemed less important. Based on the resource constraints of edge devices, the system adaptively adjusts the quantization strategy, assigning relatively higher precision to layers with greater importance and lower precision to less important layers. Additionally, when device resources allow, the system maximizes the number of high-precision layers to preserve model performance. Finally, the quantized LLMs are deployed to edge devices to handle various tasks. This intelligent adaptive quantization approach significantly reduces the storage requirements of LLMs while maintaining performance. It also enables efficient deployment of LLMs across diverse hardware platforms and various usage scenarios on the same platform. The contributions of this work can be summarized as follows:
  • A novel layer importance assessment method for LLMs is proposed, which effectively captures the relationships between semantic information across layers. This method helps to reveal the features and aspects that the model focuses on at different levels, enabling a more precise evaluation of layer importance.
  • An approach has been designed for the layer-specific quantization and dynamic deployment of LLMs based on layer importance and edge device resources, which allows for the deployment of the quantized model while striving to maintain the model’s performance.
  • LSAQ surpasses existing quantization methods of the same granularity in accuracy on 87.5% of zero-shot tasks and demonstrates superior average accuracy across all tasks. Additionally, in 90% of quantization scenarios, LSAQ outperforms comparable methods in terms of perplexity. These experimental results validate the effectiveness of LSAQ in preserving essential information while achieving efficient quantization.
II Related Work
II-A LLMs Quantization
Model quantization is a critical technique in the field of deep learning. It achieves significant model size reduction by converting model parameters (e.g., weights and activation values) from high-precision formats (typically 32-bit or 16-bit floating-point numbers) to lower-precision formats (e.g., 8-bit or 4-bit integers). This makes it feasible to deploy LLMs on resource-constrained edge devices.
There are two primary approaches to LLMs quantization: Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT incorporates quantization operations during the model training phase. It uses high-precision parameters for backpropagation and low-precision parameters for forward propagation, enabling the model to adapt to the changes introduced by quantization. In contrast, PTQ is applied after model training is complete. It reduces the representation precision of weights and activation values to shrink the model size. Compared to QAT, PTQ is more suitable for scenarios requiring rapid deployment in resource-constrained environments.
In recent research, PTQ has made significant advancements in reducing memory usage and improving computational efficiency for LLMs. The ZeroQuant [18] technique supports INT8 symmetric quantization and allows different quantization levels to be set for the final model, adapting to diverse model and task requirements. The GPTQ technique determines quantization parameters using calibration data, achieving the goal of reducing model size while maintaining accuracy. LLM.int8() improves quantization efficiency by handling activation outliers through mixed-precision decomposition. The SmoothQuant [19] technique introduces a mathematically equivalent per-channel scaling transformation to smooth activations and their corresponding weights across different channels, making the model more quantization-friendly. AWQ [20] optimizes the LLM quantization process by preserving significant weights and protecting prominent ones.
The development of these techniques has collectively facilitated the deployment and application of large language models in resource-constrained environments, offering new insights and methods for optimizing and deploying deep learning models.
Figure 2: The process of constructing top-$k$ token sets
II-B Layer Importance Assessment for LLMs
In the field of research on LLMs, prior studies like shortGPT [21] have revealed that certain layers within these models may perform similar or redundant operations. These layers exhibit a high degree of functional overlap in the information processing pipeline, leading to redundancy in their contributions to the model’s final output. By thoroughly analyzing the contribution of each layer to the model, we can determine the relative importance of individual layers. This analysis not only helps identify layers with minimal impact on model performance but also provides a theoretical foundation for model compression.
When exploring the importance of layers in LLMs, Men et al. [21] and Dumitru et al. [22] have proposed metrics such as Block Influence (BI) and Layer Redundancy (LR) to define layer importance. Both studies share a core mechanism based on calculating the cosine similarity between the input and output of each layer. Cosine similarity measures the angular similarity between two non-zero vectors; values closer to 1 indicate higher similarity, implying that the operations performed by the layer are more redundant or overlapping. In addition to using cosine similarity as a metric, Dumitru et al. [17] also estimates layer importance by analyzing the number of weights significantly exceeding the average value within a layer. If a layer contains many weights with absolute values much larger than the average weight of that layer, it is likely to have a greater impact on the model’s output.
By evaluating layer importance, the contribution of each layer to the overall functionality of the model can be quantitatively assessed. This approach provides a clear framework for analyzing and comparing the roles of different layers, facilitating the identification and optimization of key layers. As a result, it helps enhance the model’s efficiency and performance.
III Method
This section introduces LSAQ, a system designed for layer-specific adaptive quantization and dynamic deployment of LLMs on edge devices equipped with consumer-grade GPUs. As shown in Figure 1, the system is composed of both offline and online components. Due to variations in layer importance across different LLMs, as well as differences in GPU resources available on various edge devices or even on the same device under different conditions, the offline component must determine the most suitable quantization strategy based on these factors. During the online phase, the selected quantization strategy is applied to quantize the LLM to the corresponding bit widths. The quantized LLM is then deployed onto the target device, ensuring that the model can operate efficiently in resource-constrained environments while maintaining a high level of accuracy.
III-A Layer Importance Detection Module
The Layer Importance Detection (LID) module identifies the key layers that have the greatest impact on model performance by analyzing the structure and behavior of the model. This ensures that essential functions of the model are preserved during quantization.
To assess layer importance, we utilize intrinsic model features as metrics. We analyze each layer of the LLMs, extracting the hidden states at the last time step of each layer, denoted as $X_{i,in}$ and $X_{i,out}$ These hidden states capture the encoded information for the last token in the input sequence at that particular layer. Next, we multiplied these hidden states by the transpose of the embedding matrix $W_{E}^{\top}$, which maps discrete vocabulary words to a continuous vector space. This multiplication operation produces a projection of each word across the vocabulary.
Following this, we ranked the projection results to determine the indices of each word, sorted from highest to lowest based on projection values. By selecting the top-$k$ indices from the sorted results, we identified the top-$k$ most probable words for the current layer’s input and output. Finally, we converted these indices back into actual words, forming top-$k$ token sets that represent the input and output of the layer. Figure 2 illustrates the detailed process of constructing the top-$k$ token sets.
$C_{i,in}=f_{topk}({X_{i,in}\cdot W_{E}^{\top}}).$ (1)
$C_{i,out}=f_{topk}({X_{i,out}\cdot W_{E}^{\top}}).$ (2)
After obtaining the two sets $C_{i,in}$ and $C_{i,out}$, we calculate the Jaccard similarity between the two sets as a metric for assessing the importance of the layer. Jaccard similarity [23] is widely used in various fields, including machine learning and text mining, to evaluate the similarity between sets of samples or features. This similarity is calculated based on the ratio of the size of the intersection of the two sets to the size of their union. This metric is simple yet effective, providing a clear measure of the compositional similarity between the two top-$k$ token sets.
$I_{i}=1-J(C_{i,in},C_{i,out})=1-\frac{|C_{i,in}\cap C_{i,out}|}{|C_{i,in}\cup C_{i,out}|}.$ (3)
According to Equation (3), we employ the Jaccard similarity coefficient to quantify the semantic transformation degree of each layer in LLMs. Specifically, we calculate the Jaccard similarity between two top-$k$ token sets derived from the input and output of each layer and transform this value by inversion and addition to obtain the importance metric $I_{i}$ for each layer. As the Jaccard similarity coefficient increases, the importance metric $I_{i}$ decreases correspondingly.
This observation indicates that a higher similarity between the two top-$k$ token sets implies that the layer may have performed less significant semantic transformation on the input information. Consequently, it can be inferred that the layer’s contribution to semantic transformation is relatively low, leading to a lower assessment of its importance. Through Figure 3, we can visually observe the contribution of different layers to the overall performance of the Llama-2 (7B/13B) [7] and Llama-3-8B [24] models within this evaluation framework.
Figure 3: Importance of LLMs Layer
Compared to directly using cosine similarity, our approach offers a finer grained analysis by directly mapping hidden states to the vocabulary. This method captures the relationship between each layer and semantic information more effectively, helping to reveal the features and aspects each layer focuses on. As a result, it enables a more accurate evaluation of layer importance.
III-B Resource Detection Module
Effective management of GPU resources is critical for the edge-side deployment of LLMs, as it directly impacts the feasibility of deployment as well as the subsequent performance and responsiveness of the models. The Resource Detection (RD) module is designed to adapt to GPU devices of various scales and configurations, whether deployed on personal computers, servers, or cloud computing environments.
Through flexible environment variable configurations and programmatic approaches, the resource detection module can automatically select the GPU device with the most available free memory. This capability ensures efficient loading and inference of quantized LLMs, optimizing their performance during deployment.
TABLE I: Zero-shot QA task results of quantized Llama models
Model Method Avg. Bits PIQA ARC-e ARC-c BoolQ HellaSwag WinoGrande Avg. ()($\uparrow$)( ↑ )
Llama-2-7B LWQ 7 78.02 67.09 39.59 71.96 56.76 66.54 63.33
6 77.48 65.40 39.16 67.89 56.77 66.31 62.17
5 76.66 64.65 38.57 65.57 55.41 65.74 61.10
LSAQ (Ours) 7 78.02 67.09 39.59 71.96 56.76 66.54 63.33
6 77.29 65.62 39.68 69.69 56.99 65.90 62.53
5 77.15 64.74 39.19 65.99 55.55 65.59 61.37
Llama-2-13B LWQ 7 78.94 72.26 45.48 71.90 59.73 68.90 66.20
6 78.78 71.94 44.45 71.90 59.08 68.51 65.78
5 78.02 70.59 43.69 71.44 58.37 68.20 65.05
LSAQ (Ours) 7 79.00 72.73 45.73 72.51 59.78 69.22 66.50
6 78.85 72.32 45.52 71.35 59.29 68.98 66.05
5 78.56 70.88 43.26 71.28 58.55 68.82 65.23
Llama-3-8B LWQ 7 79.02 79.46 49.91 81.50 59.13 73.88 70.48
6 78.67 78.37 48.81 80.06 58.83 73.22 69.66
5 78.40 74.92 45.56 75.93 57.85 72.85 67.59
LSAQ (Ours) 7 79.28 79.76 50.77 81.77 59.67 73.56 70.80
6 78.85 78.76 49.02 80.86 59.14 73.48 70.02
5 78.63 77.02 47.27 79.94 58.24 73.48 69.10
III-C Quantization Strategy Formulation Module
The core of the Quantization Strategy Formulation (QSF) module lies in designing an optimal model quantization deployment plan based on the layer importance of the target LLM and the current GPU resource availability. This strategy aims to ensure that the quantized model minimizes accuracy loss while enabling seamless deployment and efficient inference execution.
Specifically, this module first evaluates the relationship between GPU memory availability and the memory requirements of the model during loading and inference. If the GPU memory is sufficient, the model is deployed without quantization to maintain optimal performance. If the memory is insufficient for an unquantized model but sufficient for an INT8-quantized model, the entire model undergoes INT8 quantization, as the accuracy loss from INT8 quantization is negligible.
Further, if the GPU memory is insufficient for an INT8-quantized model but can support an INT4-quantized model, the module applies INT4 quantization to layers with low importance while retaining INT8 quantization for the remaining layers. Since INT4 quantization has a greater impact on model accuracy, the strategy aims to maximize the proportion of INT8-quantized layers while ensuring that the quantized model’s memory requirements remain within the GPU’s capacity. Finally, if the GPU memory cannot accommodate even an INT4-quantized model, the quantization process will be paused until other processes release memory.
This module aims to achieve an optimal balance between model accuracy and hardware resources, ensuring adaptability to various deployment environments.
III-D Model Quantization Module
After determining the deployment strategy via the QSF module, this module performs per-channel quantization on the layers of LLMs that require quantization. Compared to the per-tensor quantization method, per-channel quantization assigns independent quantization parameters to each channel of the weight matrix. This approach allows the quantization process to more accurately capture the dynamic range of data in each channel, thereby reducing quantization errors and improving model accuracy. The quantization process can be mathematically represented as:
$W_{i}^{INTn}=round(\frac{W_{i}^{FP16}}{s_{i}})\quad s.j.t.\quad s_{i}=\frac{\max\{|W_{i}|\}}{2^{n-1}-1},$ (4)
where $W$ denotes the weight matrix of the model, $W_{i}$ represents all values in the $i$-th row of the weight matrix, and the superscript of $W$ specifies its associated data type.
To achieve quantization, we first determine the maximum absolute value of the weights in each channel, denoted as $\max\{|W_{i}|\}$. Next, based on the selected quantization bit-width $n$, a specific scaling factor $s_{i}$ is computed for each channel. This scaling factor serves to map the range of floating-point weights to the integer range required for quantization. Finally, the floating-point weights of each channel are divided by their respective scaling factors and rounded to the nearest integer. Through these steps, the weights in each channel are transformed into integer form, completing the quantization process.
IV Experiments
IV-A Experimental Setup
IV-A1 Models
To verify the feasibility of our method, we conducted experiments on some of the most popular open-source LLMs in the field of natural language processing, including Llama2-7B, Llama2-13B, and Llama3-8B, developed by Meta AI. These LLMs are based on the decoder-only Transformer architecture. Due to their extensive pretraining on large-scale language data, they demonstrate exceptional performance across a variety of NLP tasks.
IV-A2 Benchmarks
To comprehensively evaluate the performance of LLMs after quantization, we adopted a dual evaluation strategy.
Firstly, we selected six zero-shot tasks to evaluate whether the core capabilities of the models are preserved after quantization:
  • PIQA [25]: A physical interaction question-answering dataset focuses on testing the model’s ability to predict how objects interact in the physical world. The questions are typically based on everyday physical scenarios, such as the typical uses of objects or non-typical but practically feasible uses. PIQA consists of approximately 20,000 question-answer pairs, which can be multiple-choice or true/false questions.
  • ARC-e, ARC-c: These two datasets represent the easy and challenging subsets of the ARC [26] (AI2 Reasoning Challenge). ARC is a scientific reasoning dataset comprising approximately 8,000 questions and includes a 14 million scientific facts corpus to support answering these questions.
  • BoolQ [27]: A question-answering task where the model must predict whether a passage contains the answer to a specific question. These questions are typically complex and non-factual queries, requiring a ”yes” or ”no” response based on the given passage.
  • HellaSwag [28]: A dataset designed to test the commonsense reasoning ability of models. It provides a series of scenarios requiring the model to select the most appropriate outcome from multiple options. This dataset demands complex reasoning based on context rather than simple word or phrase matching.
  • WinoGrande [29]: An open-domain language understanding dataset consisting of 1.5 million natural language sentence fragments, each ranging from 4 to 14 words in length. These fragments are annotated as ”yes” or ”no” to indicate whether they align with commonsense or world knowledge.
By testing the accuracy of LLMs on the aforementioned zero-shot tasks, we can assess the reasoning and generalization capabilities of the quantized LLMs during their usage.
Additionally, we have assessed the perplexity (PPL) of quantized LLMs on the WikiText2 [30] dataset. This dataset, introduced by Merity et al., is derived from Wikipedia articles and encompasses a wealth of encyclopedic knowledge. By measuring the PPL of LLMs in language generation tasks across various levels of quantization precision, we are able to accurately evaluate the models’ predictive capabilities on the test data.
IV-A3 Baseline
In the experiment, we selected Layer-Wise Quantization (hereinafter referred to as LWQ) as the baseline method. The LWQ technique aligns with our quantization strategy in terms of granularity, as both apply differential quantization based on the importance of each layer. The LWQ method evaluates layer importance using two different metrics: one is the cosine similarity between the layer’s input and output, and the other is the number of weights in the layer that are significantly greater than the average value.
Experimental results from the LWQ paper indicate that using cosine similarity as the basis for quantization leads to better performance for the Llama series models. Based on this consideration, we decided to use cosine similarity as the layer importance metric for LWQ in formulating our quantization strategy. Since this method is independent of the underlying quantization technique, we will test it using the same quantization approach as ours.
IV-B Main Results
Tables I and III present the detailed results of our experiments. To comprehensively evaluate the application of LSAQ quantization technology to LLMs, we tested the zero-shot task accuracy and perplexity of quantized Llama-2 (7B/13B) and Llama-3 (8B) models. In these experiments, INT4 quantization was applied to 25%, 50%, and 75% of the model’s layers, while the remaining layers used INT8 quantization. This quantization scheme reduced the average bit-width of LLMs to 7 bits, 6 bits, and 5 bits, corresponding to different quantization ratios.
IV-B1 Results on Zero-shot tasks
Table I provides a detailed breakdown of the accuracy and overall average scores of various quantized models across multiple zero-shot question-answering tasks.
However, there is a special case in the experimental results. On the Llama-2-7B model, the performance of the two quantization methods is exactly the same under the condition of 7-bit quantization. This is because, although there are differences in determining the order of the first 25% least important layers between the method using cosine similarity of inputs and outputs and the method using Jaccard similarity with the vocabulary, they ultimately reach a consensus at the overall level, as shown in Table II. Consequently, both methods exhibit a high degree of consistency across all performance metrics. This consistency is reflected in both the zero-shot evaluation experiments and the subsequent perplexity experiments.
TABLE II: The First 25% Layer Numbers of Llama-2-7B under Different Layer Importance Indicators
Model Importance Indicator The First 25% Layers
Llama-2-7B Cosine 27, 26, 28, 25, 24, 29, 23, 22
Jaccard (Ours) 27, 26, 28, 24, 23, 25, 22, 29
From the experimental data, it can be observed that the LSAQ (our) quantized model generally achieves superior performance in these zero-shot tasks. Except for the 25% quantization case of Llama-2-7B (where both methods show high consistency), the LSAQ method outperforms the LWQ method in terms of average accuracy across all tasks and has an advantage in accuracy for 87.5% of individual tasks. Notably, for the ARC-e and HellaSwag tasks, our method achieves higher accuracy under a wide range of conditions. Moreover, the LSAQ method is able to maintain a high level of accuracy even at lower bit numbers.
IV-B2 Results on Perplexity
The PPL of a model is a critical metric in the field of NLP used to evaluate the performance of language models. It measures the model’s ability to predict a set of sample data, particularly in assessing the model’s predictive accuracy and generalization capability. A lower perplexity value indicates higher prediction accuracy, thereby reflecting the impact of quantization on model performance.
Table III demonstrates the impact of different quantization methods on the perplexity of three LLMs at various quantization bit widths. The experimental results indicate that, except for the condition where Llama-3-8B is quantized to 7 bits, where the LWQ quantization slightly outperforms LSAQ, in the remaining 90% of cases, the model quantized by LSAQ exhibits lower perplexity on the Wikitext2 dataset.
TABLE III: Perplexity Results of Quantized Llama Models on the Wikitext2 Dataset
Model Method 8-bits 7-bits 6-bits 5-bits 4-bits
Llama-2-7B LWQ 5.476 5.771 6.156 6.396 6.919
LSAQ 5.771 6.064 6.325
Llama-2-13B LWQ 4.886 4.958 5.074 5.200 5.403
LSAQ 4.940 5.057 5.162
Llama-3-8B LWQ 6.143 6.443 6.904 7.386 10.530
LSAQ 6.508 6.887 7.331
IV-B3 Analysis of results
By analyzing the experimental data of the quantized model on zero-shot tasks and model perplexity, it can be demonstrated that evaluating the layer importance of LLMs by constructing the top-k token sets corresponding to the input and output of each layer and calculating the Jaccard similarity of these sets is effective. This evaluation approach adeptly captures the interplay between each layer and semantic information. Consequently, it facilitates a more nuanced and accurate appraisal of the significance of each layer.
This finding reveals that compared to LWQ, the LSAQ quantization method is more effective in maintaining model accuracy, as it loses less critical information during the quantization process. Consequently, this suggests that LSAQ has potential advantages in the field of model compression and deployment on edge devices.
IV-C Quantized Model Deployment
In addition to quantization, dynamic deployment of LLMs on edge devices is also one of the core capabilities of LSAQ. To investigate whether our system meets the criteria for successful deployment of large models on edge devices, we conducted tests on the memory requirements for loading model weights of the quantized Llama-2 (7B/13B) and Llama-3 (8B) models. Besides, we examined the quantization strategies provided by the system for the Llama-2-7B model under various memory constraints to assess whether they are compliant with the requirements.
Figure 4: Memory usage of quantized model
Figure 4 clearly demonstrates the amount of memory required to load the weights of three models at different quantization precisions. When these models operate at FP16 precision, they require approximately 12.82GB, 24.36GB, and 15.14GB of memory, respectively. With the reduction of the average number of bits after quantization, the memory occupancy of all models shows a significant decrease. At their lowest, the memory requirements can be reduced to about 3.56GB, 6.79GB, and 5.18GB, respectively, which allows these models to be easily deployed on mainstream graphics cards currently available on the market.
TABLE IV: Quantization Strategies for Llama-2-7B under Varying Memory Conditions
Memory FP16 Layer INT8 Layer INT4 Layer Avg. Bits
16G 32 0 0 16
12G 0 32 0 8
8G 0 32 0 8
6G 0 32 10 6.75
4G 0 1 31 4.125
During the inference process of LLMs, intermediate computational results may be generated, which also occupy a certain amount of memory. Consequently, when devising our quantization strategy, we took into full consideration the additional demand for memory during the inference process of LLMs. Consequently, we have designed corresponding safety memory strategies based on the requirements of models of different scales.
Table IV illustrates the actual quantization strategies employed by LSAQ for deploying Llama-2-7B under various memory configurations. For instance, if a GPU has 16GB of available memory, it is entirely feasible to deploy Llama-2-7B directly on the GPU at FP16 precision without the need for quantization. However, when the GPU has only 6GB of available memory, it becomes necessary to quantize the least important 10 out of the 32 layers of Llama-2-7B to INT4 precision, while the remaining 22 layers are quantized to INT8 precision, thereby facilitating the deployment of the model.
V Conclusion
This paper introduces LSAQ, a layer-specific adaptive quantization and dynamic deployment system for large language models. LSAQ adjusts quantization strategies on hardware platforms with varying resource sizes by establishing a novel layer importance assessment mechanism, thereby significantly reducing the model’s storage requirements while maintaining model performance as much as possible. Experimental results demonstrate that the accuracy of the LSAQ-quantized model on zero-shot tasks and the perplexity of the language model are superior to existing quantization methods of the same granularit, even at lower bit widths, thus maintaining high performance. This indicates that LSAQ has potential advantages in model compression and on-device model deployment. By employing intelligent adaptive quantization methods, LSAQ can efficiently deploy quantized models in resource-constrained environments, providing a new perspective and approach for the deployment of large language models on edge devices.